2. Breakeven sales for a local manufacturing are $224924. The company manufactures and sells 2 products Basic and Expert. To breakeven it must sell 1130 Basic units and 2069 Expert units. The company’s fixed costs are $108360. The Basic product sells for $165 while the Expert unit sells for $172. The Basic unit’s $108.80 variable costs. What is the contribution margin per unit of Expert units (if necessary, round your answer to the nearest cent)?

A. $72.46

B. $30.69

c. $21.68

D. $103.15

3. If a company sells its product for $35 and has a unit cost of $14, what is the company’s markup percentage?

A.20%

B. 40%

C. 60%

D. 150%

4.Using the following unit data for a company that produces Product A and B, what is the breakeven point (round to nearest unit) when fixed costs total $143500?

Product A

Product B

Units sold last year

1845

820

Sales price

$164

$370

Cost of goods sold

119

290

Sales commission

13

30

Variable costs

132

320

Contribution margin

$ 32

$ 50


A. 993 Product A; 588 Product B

B. 2646 Product A; 1176 Product B

C. 735 Product A; 735 Product B

D. 1323 Product A; 588 Product B
Answer #1

Break even units = Fixed costs/Weighted average contribution margin per unit

(1130+2069) = 108360/Weighted average contribution margin per unit

Weighted average contribution margin per unit = $33.87

(1130*56.2 + 2069*x)/3199 = 33.87

x = $21.68 approx.

Hence, the answer is c.

3.Mark up % = (35-14)/14

= 150%

i.e. D

Weighted average contribution margin per unit = (1845*32+820*50)/2665

= $37.54 per unit

Break even units = Fixed costs/Weighted average CM Per unit

= 143500/37.54

= 3823 units

Product A = 3823*1845/2665 = 2646 units

and Product B = 1176 approx.

i.e. B